OpenStudy (anonymous):

I'll give a medal to whoever helps! Factor: 2d^2-12d+18

OpenStudy (jokersmart):

Simplifying 2d2 = 12d + -16 Reorder the terms: 2d2 = -16 + 12d Solving 2d2 = -16 + 12d Solving for variable 'd'. Reorder the terms: 16 + -12d + 2d2 = -16 + 12d + 16 + -12d Reorder the terms: 16 + -12d + 2d2 = -16 + 16 + 12d + -12d Combine like terms: -16 + 16 = 0 16 + -12d + 2d2 = 0 + 12d + -12d 16 + -12d + 2d2 = 12d + -12d Combine like terms: 12d + -12d = 0 16 + -12d + 2d2 = 0 Factor out the Greatest Common Factor (GCF), '2'. 2(8 + -6d + d2) = 0 Factor a trinomial. 2((2 + -1d)(4 + -1d)) = 0 Ignore the factor 2. Subproblem 1 Set the factor '(2 + -1d)' equal to zero and attempt to solve: Simplifying 2 + -1d = 0 Solving 2 + -1d = 0 Move all terms containing d to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -1d = 0 + -2 Combine like terms: 2 + -2 = 0 0 + -1d = 0 + -2 -1d = 0 + -2 Combine like terms: 0 + -2 = -2 -1d = -2 Divide each side by '-1'. d = 2 Simplifying d = 2 Subproblem 2 Set the factor '(4 + -1d)' equal to zero and attempt to solve: Simplifying 4 + -1d = 0 Solving 4 + -1d = 0 Move all terms containing d to the left, all other terms to the right. Add '-4' to each side of the equation. 4 + -4 + -1d = 0 + -4 Combine like terms: 4 + -4 = 0 0 + -1d = 0 + -4 -1d = 0 + -4 Combine like terms: 0 + -4 = -4 -1d = -4 Divide each side by '-1'. d = 4 Simplifying d = 4 Solution d = {2, 4}.

OpenStudy (jokersmart):

I hope THAT helps some.

OpenStudy (anonymous):

(2d-6)(d-3) if you multiply out =2d^2 -6d-3d+18 so 2d=6 so d=3 or 3